Please help me, I have been trying to find common multiples among them to maybe see if that could simplify it but I don't know what else to do, here's the question,
System of equations
11.99h + 24.99p = 245.86
h + p = 14

Question: How could you simplify the numbers used in this system to simplify the system? Would this new simplification require elimination, substitution, graphing, or making a table as a new method to solve it.

Thanks so much for your help

Answer by ReadingBoosters(3246)

I would use substitution
11.99(14-p) + 24.99p = 245.86
167.86 - 11.99p + 24.99p = 245.86
167.86 - 13p = 245.86
-13p = 245.86 - 167.86
-13p = 78
p = -6
...
The substitute the p value into either equation
h + -6 = 14
h = 14 + 6
h = 20
.....................
Have more questions?
Email: HomeworkHelpers@readingboosters.com
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.